# Zero-Knowledge Proofs in Sublinear Space

### v1

@@ 6.3 Zero-Knowledge
    If b=1b{=}1: $\pi\leftarrow\mathsf{Sim}_{\mathrm{str}}(pp,x)$.
@@
Set $\mathsf{Sim}_{\mathrm{str}}:=\mathsf{Sim}_{\mathrm{std}}$. By Theorem 4, for any fixed public randomness (or programmed random oracle), the transcript distribution (commitments/openings and auxiliary values) of $\mathsf{Prover}_{\mathrm{str}}$ matches that of $\mathsf{Prover}_{\mathrm{std}}$; in non-hiding PCSs this equality is bit-for-bit. Consequently, the Fiat–Shamir challenge inputs have the same distribution, and any distinguisher against $\Pi_{\mathrm{str}}$ yields one against $\Pi_{\mathrm{std}}$ with the same advantage.
Instantiation note. If the PCS is not hiding by default (e.g., plain KZG), we assume a hiding instantiation or the standard polynomial-masking steps of the base protocol that render the overall scheme zero-knowledge. This transformation does not alter those steps; it only reorders computation and compresses memory. ∎
@@ 7.1 Performance Comparison
The primary benefit of this construction is a square-root-space improvement: from linear space $\Theta(T)$ down to $O(\sqrt{T})$ (Section 5), i.e., a factor of $\approx\sqrt{T}$ less memory. [^1] Throughout, lower-order polylogarithmic terms are suppressed; see Theorem 5. To illustrate the practical impact, we compare the streaming prover against a standard linear-space prover for a ZKP system that materializes the full execution trace in memory.
@@
Prover Space $\Theta(T)$ $O(\sqrt{T})$ (up to $\log\log$)
Prover Time $O(T\log N)$ or $O(T\log T)\dagger$ $O(T\log N)\cdot\mathrm{poly}(\lambda)$
@@
$\dagger$ Prover time is $O(T\log N)$ for KZG/IPA, or $O(T\log T)$ for STARK/FRI.
$\ddagger$ The “identical” entries hold for linear PCS instantiations (e.g., KZG/IPA), by Theorem 4. For hash-based STARK/FRI commitments, the prover remains sublinear-space but transcripts (and thus some constants) generally differ.
This construction dramatically improves the prover’s memory footprint while keeping time complexity comparable up to polylogarithmic factors. Crucially, the proof itself and the verifier’s work remain unchanged for linear PCS instantiations, ensuring that the prover can be a drop-in replacement in existing ZKP ecosystems without affecting end-users or protocol security parameters.
@@ 7.2 Concrete Impact on System Requirements
To contextualize the magnitude of the space improvement, consider two representative computation sizes. We assume each trace element occupies 32 bytes (e.g., a 256-bit field element), and suppress constant factors from the number of registers $k=O(1)$ and small auxiliary state. With $\lambda=\Theta(\log T)$, the dominant space bound is $O(\sqrt{T})$ up to a $\log\log$ factor (Section 5).
@@
  • Streaming prover: $\sqrt{2^{24}}\!\times 32=2^{12}\!\times 32=131{,}072$ bytes $\approx\mathbf{0.13}$ MB.
@@
  • Streaming prover: $\sqrt{2^{30}}\!\times 32=2^{15}\!\times 32=1{,}048{,}576$ bytes $\approx\mathbf{1.0}$ MB.
@@ 7.4 Limitations and Future Work
This work presents both a significant advance and several exciting avenues for future research.
  • Tightening Space Bounds: The bound is $O(\sqrt{T}\cdot\log\log T)$ space (equivalently, $O(\sqrt{T})$ up to a $\log\log$ factor). It is natural to ask whether constants and lower-order terms can be further reduced, or whether matching lower bounds can be proved under general read-degree $r_{\text{reg}}$ and register count $k$.
  • Time Complexity Optimization: While the polylogarithmic overhead is acceptable in many memory-bound scenarios, reducing runtime remains valuable. Custom tree-evaluation routines tailored to AIR/PCS structure (e.g., avoiding generic interpolation where possible) could lower the $\log$ factors.
  • Hash-/lattice-based PCSs (STARK/FRI) and PQ settings: The identical-transcript guarantee relies on linear PCS commits (KZG/IPA). With Merkle/FRI-style hash commitments one can still stream with $o(T)$ space by carrying Merkle paths/FRI layers in $\mathsf{aux}$, but the resulting transcripts generally differ (different leaf chunking and digests). This raises $b_{\text{val}}$ (and thus the $\Delta\!\cdot\!b_{\text{val}}$ term in Cook–Mertz space). Quantifying the exact overhead while retaining sublinear space, and characterizing when transcript equivalence can be restored, are important open problems for post-quantum instantiations.
@@ 8 Conclusion
We have presented, to our knowledge, the first zero-knowledge proof system with sublinear prover space, breaking a long-standing barrier in verifiable computation. The core technical contribution behind this work is the Tree Evaluation Equivalence (Theorem 4), which establishes a mathematical equivalence between cryptographic proof generation and the Tree Evaluation problem. By combining the structural decomposition of Williams with a PCS whose commit map is linear, we express ZKP commitment generation as a recursive function $F_{v}$ over a tree, then evaluate it in low space using the Cook–Mertz algorithm. This yields a space improvement from $\Theta(T)$ to $O(\sqrt{T})$ up to a $\log\log$ factor.
@@
This construction preserves the security properties, proof size, and verification cost of existing linear-space systems, allowing it to serve as a direct, space-efficient replacement in a wide variety of applications (for linear PCS instantiations). Beyond the immediate practical impact, this work suggests a broader pattern of algorithmic transference, wherein advances in space-efficient algorithms for abstract computational models can be systematically lifted to create space-efficient cryptographic protocols for those same models. This observation points to a rich landscape where breakthroughs in algorithm design directly fuel improvements in cryptographic protocol efficiency and practical deployment.

### v2

@@ 6.3 Zero-Knowledge
    If b=1b{=}1: $\pi\leftarrow\mathsf{Sim}_{\mathrm{std}}(pp,x)$.
@@
Set $\mathsf{Sim}_{\mathrm{str}}:=\mathsf{Sim}_{\mathrm{std}}$. The same simulator for the baseline works because the distribution of FS inputs and outputs is identical (bit-identical for non-hiding PCSs; distribution-identical for hiding PCSs).
By Theorem 4, for any fixed public randomness (or programmed random oracle), the transcript distribution (aggregate commitments and openings) of $\mathsf{Prover}_{\mathrm{str}}$ matches that of $\mathsf{Prover}_{\mathrm{std}}$; in non-hiding PCSs this equality is bit-for-bit. In hiding PCSs, the per-block blinders sum to a uniform blinder per polynomial (Lemma 2), so commitment/opening distributions are unchanged. Consequently, the Fiat–Shamir challenge inputs have the same distribution, and any distinguisher against $\Pi_{\mathrm{str}}$ yields one against $\Pi_{\mathrm{std}}$ with the same advantage.
Instantiation note. If the PCS is not hiding by default (e.g., plain KZG), we assume a hiding instantiation or the standard polynomial-masking steps of the base protocol that render the overall scheme zero-knowledge. Our transformation merely reorders computation and compresses memory; it does not alter those steps. ∎
Remark (non-linear PCSs).
For Merkle/FRI-style PCSs, the streaming prover remains sublinear-space, but transcripts generally differ (e.g., chunking and digest structure). Security still follows from the underlying commitment binding/collision resistance and the unchanged IOP analysis; our prover is a black-box reimplementation preserving those guarantees without claiming transcript identity. Note that $b_{\text{val}}$ may increase but still yields $o(T)$ space.
@@ 7.1 Performance Comparison
The primary benefit of this construction is a square-root-space improvement: from linear space $\Theta(T)$ down to
$O\!\big(\sqrt{T}+\log T\log\log T\big)\;=\;O\!\big(\sqrt{T}\cdot\mathrm{polylog}\,T\big)$
(Section 5), i.e., a factor of $\approx\sqrt{T}$ less memory up to lower-order polylogarithmic terms. [^1] Throughout, lower-order polylogarithmic terms are made explicit via Theorem 5 and the Cook–Mertz stack analysis. To illustrate the practical impact, we compare the streaming prover against a standard linear-space prover for a ZKP system that materializes the full execution trace in memory.
@@
Prover Space $\Theta(T)$ $O\!\big(\sqrt{T}+\log T\log\log T\big)\ast$
Prover Time $O(T\log N)$ or $O(T\log T)\dagger$ $O(T\log N)\cdot\mathrm{poly}(\lambda)$$\S$
@@
$\ast$ Space bound from Theorem 5; blocking by $b_{\text{blk}}$ also improves cache locality.
$\dagger$ Prover time is $O(T\log N)$ for KZG/IPA (evaluation-basis MSMs) or $O(T\log T)$ for STARK/FRI.
$\S$ From Theorem 5: constant number of streaming passes plus polylog traversal stack.
$\ddagger$ ”Identical” holds for linear PCSs (e.g., KZG/IPA) when (i) the same basis/SRS and encoding are used, and (ii) only the aggregate per-polynomial commitments (not per-block commitments) are hashed into Fiat–Shamir, in the same order as the baseline. For hash-based STARK/FRI commitments, the prover remains sublinear-space but transcripts (and some constants) generally differ.
This construction dramatically improves the prover’s memory footprint while keeping time complexity comparable up to polylogarithmic factors. For linear PCS instantiations with the aggregate-only-into-FS discipline, existing verifier code and parameters remain completely unchanged, enabling drop-in replacement without affecting end-users or protocol security parameters.
@@ 7.2 Concrete Impact on System Requirements
To contextualize the magnitude of the space improvement, consider two representative computation sizes. We assume each trace element occupies 32 bytes (e.g., a 256-bit field element), and the true bound is $O\!\big(k\cdot\sqrt{T}+\log T\log\log T\big)$ with $k=O(1)$ registers. With $\lambda=\Theta(\log T)$, the dominant space bound is $O\!\big(\sqrt{T}+\log T\log\log T\big)$ (Section 5).
@@
  • Streaming prover: $\sqrt{2^{24}}\!\times 32=2^{12}\!\times 32=131{,}072$ bytes $\approx\mathbf{0.13}$ MB + a small polylog overhead (well below $1$ MB at this scale).
@@
  • Streaming prover: $\sqrt{2^{30}}\!\times 32=2^{15}\!\times 32=1{,}048{,}576$ bytes $\approx\mathbf{1.0}$ MB + a small polylog overhead (still $\ll$ the $\sqrt{T}$ term).
@@ 7.4 Limitations and Future Work
This work presents both a significant advance and several exciting avenues for future research. We emphasize three critical caveats:
  • Transcript identity requirements: Identical transcripts require linear PCS + basis/SRS match + aggregate-only Fiat–Shamir. Violating these conditions changes challenges and thus transcripts.
  • Hash-based PCS limitations: For STARK/FRI-style commitments, transcripts generally differ due to chunking and digest structure, though sublinear space is retained.
  • Constant factors at small scale: For small $T$, constants, I/O costs, or setup overhead may dominate the asymptotic $O(\sqrt{T})$ improvement.
Additional research directions include:
  • Tightening Space Bounds: Can constants and lower-order terms be further reduced? Can we achieve the log-space tree evaluation goal? Are matching lower bounds possible under general read-degree $r_{\text{reg}}$ and register count $k$?
  • Time Complexity Optimization: Custom tree-evaluation routines tailored to AIR/PCS structure could lower the polylog factors.
  • Post-Quantum Extensions: Developing PQ-friendly transcript-matching techniques for hash-based commitments while quantifying the exact overhead in $b_{\text{val}}$ and maintaining sublinear space.
@@ 8 Conclusion
We have presented, to our knowledge, the first zero-knowledge proof system with sublinear prover space, breaking a long-standing barrier in verifiable computation. The core technical contribution behind this work is the Tree Evaluation Equivalence (Theorem 4), which establishes a mathematical equivalence between cryptographic proof generation and the Tree Evaluation problem. By combining the structural decomposition of Williams with a PCS whose commit map is linear, we express ZKP commitment generation as a recursive function $F_{v}$ over a tree, then evaluate it in low space using the Cook–Mertz algorithm. This yields a space improvement from $\Theta(T)$ to
$O\!\big(\sqrt{T}+\log T\log\log T\big)\;=\;O\!\big(\sqrt{T}\cdot\mathrm{polylog}\,T\big).$
@@
For linear PCS instantiations (e.g., KZG/IPA) and when the same basis/SRS and encodings are used, we further enforce that only the aggregate per-polynomial commitments (not per-block commitments) are hashed into Fiat–Shamir in the same order as the baseline; under these conditions, the proof transcript distribution is identical (bit-for-bit in non-hiding variants). For STARK/FRI-style hash commitments, the prover remains sublinear-space but transcripts typically differ.
Finally, the prover executes a constant number of streaming passes (wires, permutation, quotient, openings), each linear in $T$, so time remains $O(T\cdot\mathrm{polylog}\,N)$ for KZG/IPA (and $O(T\log T)$ for FRI/STARK), while the Cook–Mertz traversal contributes only a polylogarithmic additive stack term due to the constant arity of our computation tree.
@@
Code Availability. A reference implementation of the sublinear-space zero-knowledge proof system described in this work is available at: